OpenStudy (anonymous):

the reciprocal of three plus the reciprocal of 6 is the reciprocal of what number please show steps thank u :)

OpenStudy (saifoo.khan):

\[\frac{1}{3} + \frac{1}{6} \] LCM \[\frac{3}{6} \] \[\frac{1}{2}\]